OpenStudy (anonymous):

I have 7z - 3x = 18 2z = 1 + 5x And now I dk what to do

OpenStudy (anonymous):

ok thanks so much

OpenStudy (anonymous):

x=1 and z=3

OpenStudy (anonymous):

The is a literal equation with substitution I flipped the 2z = 1 + 5x It was originally 2z - 5x = 1

OpenStudy (anonymous):

it doesnt matter:)

OpenStudy (amistre64):

divide by 2 to get z alone

OpenStudy (anonymous):

I divide what by two?

OpenStudy (amistre64):

2z = 1 + 5x -----------; z = (1/2) +(5/2)x 2

OpenStudy (anonymous):

What about the first equation?

OpenStudy (amistre64):

what about it?

OpenStudy (anonymous):

What happens to it

OpenStudy (amistre64):

you sub the z value into it and solve for x

OpenStudy (amistre64):

7z - 3x = 18 7((1/2)+(5/2)x) - 3x = 18

OpenStudy (anonymous):

Then what do I do?

OpenStudy (amistre64):

7((1/2)+(5/2)x) - 3x = 18 (2) 7(1+5x) - 6x = 36

OpenStudy (amistre64):

7+35x -6x = 36 x = 1

OpenStudy (anonymous):

Ugh. Now I should have known that!

OpenStudy (amistre64):

i goofed that answer but you see it lol

OpenStudy (anonymous):

What about the 7?

OpenStudy (anonymous):

Doesn't it get multiplied by 2 also?

OpenStudy (amistre64):

7+35x -6x = 36 29x = 29 ..... x=1; lol thats a coincidents

OpenStudy (amistre64):

lets focus on that part and see what I did ok?

OpenStudy (amistre64):

1+5x 7( ----- ) times 2 = 7(1+5x) 2

OpenStudy (anonymous):

How do I find the Y again?

OpenStudy (anonymous):

Because I am confused about that bit right know?

OpenStudy (amistre64):

the y or the z?

OpenStudy (anonymous):

Sorry, Z

OpenStudy (amistre64):

ok... do you see why when we mulitply by 2 we get 7(1+5x) tho?

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

then lets solve for x: 7+35x -6x = 36 ; -7 and add like terms 7-7+35x -6x = 36-7 29x = 29 x = 1 right?

OpenStudy (anonymous):

Yes I have the much of the problem solved :D

OpenStudy (amistre64):

2z = 1 + 5x 2z = 1+5 2z = 6 z = 3

OpenStudy (anonymous):

Finally the problem is finished :D

OpenStudy (amistre64):

once you know what x is....use it lol

OpenStudy (anonymous):

Thks again!!!!!!!!!!!!!!

OpenStudy (amistre64):

yw :)

OpenStudy (anonymous):

The problem looks like this 7z - 3x = 18 2z - 5x = 1 changes to 2z = 1 + 5x Then I divide 2z by 2 to get 7z - 3x = 18 z = 1/2 5x/2 which turns to 7(1/2) 5x/2) - 3x = 18 which turns to 7 + 35x -6x = 36 which turns to 7 + 29x = 36 which turns to 29x = 29 then x = 1 then 2z = 1 + 5 (1) 2z = 1 + 5 2z = 6 z = 3 (1,3)